OpenStudy (johnweldon1993):

Differential Equations, Mixing problem. A tank contains 1160 L of pure water. A solution that contains 0.03 kg of sugar per liter enters a tank at the rate 7 L/min The solution is mixed and drains from the tank at the same rate. a.) How much sugar is in the tank initially? b.) Find the amount of sugar in the tank after t minutes. c.) Find the concentration of sugar in the solution in the tank after 72 minutes.

OpenStudy (johnweldon1993):

I made s(t) be the amount of salt at time 't' ds/dt = rate in - rate out Well the rate in would be the concentration times the rate... so ds/dt = (.03kg)(7L/min) - (s/1160)(7L/min) ds/dt = .21 - (7s/1160) Kind of get confused from there...I figure I'll have to get the 's' on one side right? separable? I wanted to let 7/1160 = a...so I would just have ds/dt +as = .21 Am I even on the right track here? Just starting out in the class.

OpenStudy (austinl):

I am a bit rusty, but try these notes. http://tutorial.math.lamar.edu/Classes/DE/Modeling.aspx

OpenStudy (johnweldon1993):

Hmm I don't see much that explains what the next step would be...but thanks for the link!

OpenStudy (johnweldon1993):

This is a linear Diff. eq correct?

OpenStudy (austinl):

I believe so.

OpenStudy (johnweldon1993):

Hmm I'm going to try and follow those notes then... Integration Factor \[\large u(t) = e^{\frac{7}{1160}t}\] Multiply everything by that \[\large e^{\frac{7}{1160}t} \frac{ds}{dt} + \frac{7}{1160}e^{\frac{7}{1160}t}s = .21e^{\frac{7}{1160}t}\] Simplify \[\large (e^{\frac{7}{1160}t}s)^t = .21e^{\frac{7}{1160}t}\] Integrate both sides \[\large e^{\frac{7t}{1160}} s + k = 34.8e^{\frac{7t}{1160}} + c\] Constants on 1 side \[\large e^{\frac{7t}{1160}} s= 34.8e^{\frac{7t}{1160}} + c\] Divide by \[\large e^{\frac{7t}{1160}}\] \[\large s(t) = 34.8 + ce^{\frac{-7t}{1160}}\] Hows that look? lol

OpenStudy (johnweldon1993):

Since we know then that s(0)- 0 we can solve for c 0 = 34.8 + c(1) so C = 34.8 so my final equation would be \[\large s(t) = 34.8 + 34.8e^{\frac{-7t}{1160}} \] Hows this look?

OpenStudy (johnweldon1993):

And apparently that isn't correct...sigh

OpenStudy (experimentx):

seems okay

OpenStudy (experimentx):

as t->infinity ... the solution must converge to input rate. and at t=0, the solution must be in it's initial condition. just note these to know if you are correct or not.

OpenStudy (johnweldon1993):

Well i guess that makes sense it isnt correct...when I plug t = 0 into here I don't get 0

OpenStudy (johnweldon1993):

yeah and this approaches infinity if I'm correct...so I guess start over lol

OpenStudy (experimentx):

did you apply initial condition?

OpenStudy (johnweldon1993):

Well yeah, isnt tht what you do to solved for 'c' ? Since s(0) = 0 ...plugging that into the formula I have give you 0 = 34.8 + c Which means c would = -34.8 Omg -_- hang on that's it isnt it...I dropped that negative sign before

OpenStudy (experimentx):

woops!! something went wrong with MMA.

OpenStudy (johnweldon1993):

Well then at 72 minutes that would give me .3 which is incorrect

OpenStudy (experimentx):

\[ \large s(t) = 34.8 + ce^{\frac{-7t}{1160}} \\ 0 = 34.8 + ce^{0} \implies -34.8 = c\] rest is okay.

OpenStudy (experimentx):

probably i made miscalculation on concentration ... rest is okay.

OpenStudy (johnweldon1993):

Yeah I finally got the equation down...just now at 72 minutes my equation is messing me up I get 12.26 ...that seems correct to me...but says incorrect online...sigh I'm going to hate this class lol

OpenStudy (johnweldon1993):

Wait a minute...I finding the concentration of sugar in the tank after 72 minutes, Is that different than the AMOUNT of sugar in the tank after 72 minutes?

OpenStudy (experimentx):

amount of sugar = volume times concentration ... of course.

OpenStudy (johnweldon1993):

oh crap that's right...lol thank you for walking me through dumb mistakes like that @experimentX

OpenStudy (experimentx):

yw

OpenStudy (johnweldon1993):

And that's correct! Thank's again!